consider the chemical reaction between carbon and oxygen to produce carbon dioxide. what mass of carbon dioxide will be produced if 32.24 grams of carbon reacts completely with 54.75 grams of oxygen?
Answer #1

The reaction that takes place is

C + O2 => CO2

1 mole of Carbon reacts completely with 1 mole of Oxygen and produce 1 mole of CO2

given mass of carbon 32.24 grams

moles of carbon = 32.24 / 12 (molecular mass of Carbon) = 2.69

given mass of oxygen 54.75 grams

moles of O2 = 54.75 / 32 (molecular mass of O2) = 1.711

So limiting reagent in this reaction is O2

so moles of CO2 produced is equal to moles of O2, as 1 mole of O2 produce 1 mole of CO2

so moles of CO2 = 1.711

mass = 1.711 * 44 (molecular mass of CO2) = 75.28 grams

mass of carbon dioxide produced is 75.28 grams